OpenStudy (anonymous):

inverse laplace transform of (-2e^-2s)(s-4)/(s-5)^3

OpenStudy (anonymous):

is it \[\frac{ -2e^{-2s}*(s-4) }{ (s-5)^3 }\]

OpenStudy (anonymous):

yes, it is.

OpenStudy (anonymous):

according to matlab -e^(5t-10)*(t-2)*t*theta(t-2) but not sure, i havent dont it for a while

OpenStudy (anonymous):

sorry..i didn't use matlab..by the way how can i do partial fraction for the dinominator??

OpenStudy (sirm3d):

for the rational expression \[\frac{ s-4 }{ (s-5)^3 }=\frac{ A }{ (s-5) }+\frac{ B }{ (s-5)^2 }+\frac{ C }{ (s-5)^3 }\]

OpenStudy (anonymous):

tq very much...that's what i'm looking for.. =D

OpenStudy (sirm3d):

yw